OpenStudy (erniellyd28):

a laboratry technician needs to make a 60 liter batch of a 40% acid solution. how can the laboratory technician combine a batch of an acid solution that is pure acid with another that is 20% to get the desired concentration?

OpenStudy (peachpi):

40% of 60 liters = 0.4*60 = 24 liters of acid needed 60 - 24 = 36 liters of the other liquid All the other liquid has to come from the 20% solution. 80% of x liters = 36 liters x = 45 liters So you need 45 liters of the 20% solution and 15 liters of the pure acid.

OpenStudy (erniellyd28):

thank you very very much :-)

OpenStudy (erniellyd28):

if i could give you more medals i would :-)